OpenStudy (anonymous):

Rationalize the denominator and simplify sqrt7/sqrt3 is the answer sqrt21?

OpenStudy (anonymous):

(sqrt21)/3

OpenStudy (ash2326):

We have \[\frac{\sqrt{7}}{\sqrt{3}}\] Multiply numerator and denominator by \(\sqrt{3}\)\[\frac{\sqrt{7}\times \sqrt 3}{\sqrt{3}\times \sqrt3}\] we get \[\frac{\sqrt{21}}{3}\]

OpenStudy (anonymous):

Thank-you

OpenStudy (ash2326):

:D

OpenStudy (anonymous):

hey, @ash2326 , how u write the fraction line?

OpenStudy (ash2326):

@arnab come to maths chat. I'll tell you

OpenStudy (anonymous):

ok